What is the answer for 8 to the x power = 16? Solve for x.
Answer by Boreal(15235)

8^x=16
2^3x=16, since 2^3=8
16 is 2^4
therefore 2^3x=2^4
3x=4
x=(4/3)
8^(4/3)= cube root of 8 (2) raised to the fourth power (16).